# Vision Language Models See What You Want but not What You See

### v5

@@ 1 Introduction
Intentionality is the capacity of the mind to be directed toward, represent, or stand for objects, properties, or states of affairs for further executable actions (Anscombe, 1956). To say one could understand intentionality is to say one has the capacity to comprehend the mental content for action in another mind (Premack & Woodruff, 1978; Rosenthal, 1991). This capacity has been seen as a key distinction between humans and machines (Searle, 1980). It is argued that despite well manipulation of language symbols, machines cannot understand intentional meanings of others because it lacks theory-of-mind (ToM), the kind of abilities that allows one to infer the mental content of others (Premack & Woodruff, 1978; Dennett, 1987). Nevertheless, several recent studies have showed that large language models (LLMs) and their supporting vision language models (VLMs) exhibit ToM abilities (Kosinski, 2023; Strachan et al., 2024; Shi et al., 2024), thus calling for updated examinations of the nature of ToM and the potential for current and future artificial intelligence to possess such abilities.
@@
We believe an important approach to said inquiry is examining the extent to which different ToM abilities necessitate model-based reasoning. Specifically, a distinction can be drawn between ToM abilities based on simulation-theory and theory-theory. The former involves the construction of an internal model of self-other relations to reason about the mental states of others, whereas the latter requires only the use of theoretical knowledge regarding the relations between mind and behavior (Gopnik & Wellman, 1992; Frith & Frith, 2005; Shanton & Goldman, 2010). Whether current artificial intelligence systems possess internal models that are available for reasoning remains a key debate, with several influential accounts questioning the existence of model-based reasoning among LLMs (Hao et al., 2023; Mitchell & Krakauer, 2023; Yildirim & Paul, 2024; Goddu et al., 2024). If this is indeed the case, then evidences regarding the possession of ToM abilities in VLMs above would imply that ToM abilities do not require mental simulation, and that mental simulation is not within the foundational capabilities of ToM systems.
@@
We tested this critical prediction by assessing VLMs’ ability to perform intentionality understanding and level-2 perspective-taking. ToM is commonly understood to be grounded in perspective-taking, a series of multi-level abilities that involves the cognitively undertaking of the perspective of another (Barnes-Holmes et al., 2004). Level-1 perspective-taking refers to the acknowledgement that different people can see different things, whereas level-2 perspective-taking involves the understanding of how another person may see the same thing differently. While level-1 perspective-taking emerges in humans as early as 2 years old, much older children are found to struggle with level-2 perspective-taking (Piaget, 1977). This is likely because despite its relatively low level in the perspective-taking hierarchy, this ability requires model-based reasoning, exemplified in the visual domain as inferences based on mental rotation (Lehmann & Jansen, 2019; Gunia et al., 2021). On the other hand, while intentionality understanding involves high-level cognition and abstract reasoning, it is unclear whether this complex ability necessitate mental simulations (Kilner, 2004; Bianco et al., 2024). Assessing level-2 perspective-taking and intentionality understanding in VLMs could thus provide insights into not only VLMs’ abilities within these two ToM domains but also how they mechanistically related in terms of information-processing. However, these areas remain largely unexplored in the current literature, and in intelligent systems beyond human beings.
@@
To address this critical gap in the literature, we introduced IntentBench and PerspectBench, two targeted benchmarks designed to systematically evaluate the capabilities of current VLMs in intentionality understanding and level-2 perspective-taking, respectively.
@@ 2.2.1 Level-2 Perspective-taking
In Piagetian developmental psychology, the acquisition of level-2 perspective-taking ability marks a milestone of human cognitive development for that it indicates the elimination of egocentrism – the inability to consider perspectives other than one’s own (Piaget, 1977). The Three Mountain Task invented by Piaget is widely used in developmental psychology laboratories as the gold standard for testing level-1 and level-2 perspective-taking abilities in children (Piaget & Inhelder, 1957; Johnson, 1975; Ford, 1979; Lamm et al., 2007). In a standard Three Mountain Task assessment, a child is instructed to position oneself in front of a model featuring three mountains. These mountains vary in size and are distinguished by unique characteristics: one is covered in snow, another has a red cross at its peak, and the third is topped with a hut. The child is then asked to perform a complete 360-degree examination of the model. Subsequently, another individual is introduced and takes a different vantage point to observe the model. The child is presented with several photographs that showcase various viewpoints of the model and is tasked with identifying which photograph accurately represents what the other person sees. At around four years of age, children typically select the photograph that matches their own perspective. By six years old, they begin to acknowledge viewpoints that differ from their own, and by the ages of seven to eight, they are generally able to reliably identify the perspective of another individual (Piaget & Inhelder, 1969). To test level-1 and level-2 perspective-taking in VLMs, we adapted the Three Mountain task into formats that are suitable for benchmarks with minimal confounding details while preserving real-life spatiality. In particular, we used groups of 3-4 commonly-seen elastic cans organized into different spatial patterns to mimic the mountain model. Like in the original task, we used a doll placed to face the organization from different angels as the object of perspective-taking.
@@ 2.2.2 Intentionality Understanding
2.3 Model Selection and Experiment
We evaluated three categories of VLMs. To ensure a fair comparison, all VLMs are evaluated on their ability to reason over images and texts under a zero-shot generation task. A complete list of models is reported in the results section as shown in Figure 3. Model size data are curated at the same time. The models are categorized as follows:
@@ 2.3 Model Selection and Experiment
    Open-source VLMs with Multi-Image Reasoning: Includes models with different sizes and other variants such as CogVLM Series (Hong et al., 2024), Qwen series(Qwen-VL (Bai et al., 2023), Qwen-2 (Wang et al., 2024)), and Blip2 (Li et al., 2023), LLaVA-Next (Liu et al., 2024) , which are capable of reasoning over interleaved multiple images and texts.
@@
In total, we processed 37 models for evaluation. All the model performances in intentionality understanding and perspective-taking together with human baseline performances are presented here (Figure 2). In order to analyze the reasoning abilities of VLMs, we ask the models to explain their answers after they have given the answers.
@@ 3.1 Overall Performance
Figure 4: VLMs perform significantly better in intentionality understanding compared to perspective-taking. Paired samples t-test: $p=2.62\times 10^{-19}$, $t=17.651$
@@ 3.2 Relationship Between Model Performance and Model Size
To quantitatively assess the impact of model size on performance, we conducted a linear regression analysis of accuracy scores against model size across our 37 VLMs. For intentionality understanding, the coefficient of determination is $R^{2}=0.2797$, with a slope of $0.0599$ and an offset of $0.3925$. This positive slope indicates that as model size increases, intentionality understanding improves, which aligns with the scaling law hypothesis—the expectation that larger models generally exhibit better performance on cognitive tasks. The results for perspective-taking reveal a strikingly different trend. The coefficient of determination is much lower at $R^{2}=0.0176$, with a slope of $-0.0057$ and an offset of $0.1437$. The negative slope suggests that, contrary to expectations, larger models do not show significant improvements in perspective-taking ability. In fact, their performance remains largely unchanged or even slightly decreases as they scale. This finding directly contradicts the scaling law hypothesis and suggests that perspective-taking may require fundamentally different cognitive processes that are not naturally enhanced through increased model size alone. This dissociation between intentionality understanding and perspective-taking as models scale highlights a potential limitation in current VLMs. While intentionality understanding appears to benefit from increased parameters and training data, perspective-taking does not seem to follow the same trajectory.

### v6

@@ 1 Introduction
Intentionality is the capacity of the mind to be directed toward, represent, or stand for objects, properties, or states of affairs for further executable actions (Anscombe, 1956). To say one could understand intentionality is to say one has the capacity to comprehend the mental content for action in another mind (Premack & Woodruff, 1978; Rosenthal, 1991). This capacity has been seen as a key distinction between humans and machines (Searle, 1980). It is argued that despite well manipulation of language symbols, machines cannot understand intentional meanings of others because it lacks theory-of-mind (ToM), the kind of abilities that allows one to infer the mental content of others (Premack & Woodruff, 1978; Dennett, 1987). Nevertheless, several recent studies have showed that large language models (LLMs) and their supporting vision language models (VLMs) exhibit ToM abilities (Kosinski, 2023; Li et al., 2025; Strachan et al., 2024; Shi et al., 2024), thus calling for updated examinations of the nature of ToM and the potential for current and future artificial intelligence to possess such abilities.
@@
We believe an important approach to said inquiry is examining the extent to which different ToM abilities necessitate model-based reasoning. Specifically, a distinction can be drawn between ToM abilities based on simulation-theory and theory-theory. The former involves the construction of an internal model of self-other relations to reason about the mental states of others, whereas the latter requires only the use of theoretical knowledge regarding the relations between mind and behavior (Gopnik & Wellman, 1992; Frith & Frith, 2005; Shanton & Goldman, 2010). Whether current artificial intelligence systems possess internal models that are available for reasoning remains a key debate, with several influential accounts questioning the existence of model-based reasoning among LLMs (Hao et al., 2023; Mitchell & Krakauer, 2023; Yildirim & Paul, 2024; Li et al., 2022; Goddu et al., 2024). If this is indeed the case, then evidences regarding the possession of ToM abilities in VLMs above would imply that ToM abilities do not require mental simulation, and that mental simulation is not within the foundational capabilities of ToM systems.
@@
We tested this critical prediction by assessing VLMs’ ability to perform intentionality understanding and level-2 perspective-taking. ToM is commonly understood to be grounded in perspective-taking, a series of multi-level abilities that involves the cognitively undertaking of the perspective of another (Barnes-Holmes et al., 2004). Level-1 perspective-taking refers to the acknowledgement that different people can see different things, whereas level-2 perspective-taking involves the understanding of how another person may see the same thing differently. While level-1 perspective-taking emerges in humans as early as 2 years old, much older children are found to struggle with level-2 perspective-taking (Piaget, 1977). This is likely because, despite its relatively low level in the perspective-taking hierarchy, this ability requires model-based reasoning, exemplified in the visual domain as inferences based on mental rotation (Lehmann & Jansen, 2019; Gunia et al., 2021). On the other hand, as another ability at the core of human Theory-of-Mind, while intentionality understanding involves high-level cognition and abstract reasoning, it is unclear whether this complex ability necessitate mental simulations(Wellman, 1992; Frith & Frith, 2006; Apperly, 2010; Kilner, 2004; Bianco et al., 2024). Assessing level-2 perspective-taking and intentionality understanding in VLMs could thus provide insights into not only VLMs’ abilities within these two ToM domains but also whether the comparative performance between them follows that of humans. However, these areas remain largely unexplored in the current literature, and in intelligent systems beyond human beings.
@@
To address this critical gap in the literature, we leveraged and further instantiated the IntentBench and PerspectBench of CoreCognition benchmark (Li et al., 2024b), two targeted dataset designed to systematically evaluate the capabilities of current VLMs in intentionality understanding and level-2 perspective-taking, respectively.
@@ 2.2.1 Level-2 Perspective-taking
In Piagetian developmental psychology, the acquisition of level-2 perspective-taking ability marks a milestone of human cognitive development as it indicates the elimination of egocentrism – the inability to consider perspectives other than one’s own (Piaget, 1977). The Three Mountain Task invented by Piaget is widely used in developmental psychology laboratories as the gold standard for testing level-1 and level-2 perspective-taking abilities in children (Piaget & Inhelder, 1957; Johnson, 1975; Ford, 1979; Lamm et al., 2007). In a standard Three Mountain Task assessment, a child is instructed to position oneself in front of a model featuring three mountains. These mountains vary in size and are distinguished by unique characteristics: one is covered in snow, another has a red cross at its peak, and the third is topped with a hut. The child is then asked to perform a complete 360-degree examination of the model. Subsequently, another individual is introduced and takes a different vantage point to observe the model. The child is presented with several photographs that showcase various viewpoints of the model and is tasked with identifying which photograph accurately represents what the other person sees. At around four years of age, children typically select the photograph that matches their own perspective. By six years old, they begin to acknowledge viewpoints that differ from their own, and by the ages of seven to eight, they are generally able to reliably identify the perspective of another individual (Piaget & Inhelder, 1969). To test level-2 perspective-taking in VLMs, we adapted the Three Mountain task into formats that are suitable for benchmarks with minimal confounding details while preserving real-life spatiality. In particular, we used groups of 3-4 commonly-seen elastic cans organized into different spatial patterns to mimic the mountain model. Like in the original task, we used a doll placed to face the organization from different angels as the object of perspective-taking.
@@ 2.2.2 Intentionality Understanding
2.3 Examined Vision Language Models
Recent advances in multi-modal learning have been largely driven by the unified modeling of visual and textual modalities using transformers (Li et al., 2019; Xu et al., 2023; Tan & Bansal, 2019; Alayrac et al., 2022; Radford et al., 2021). With the rise of large language models (LLMs), state-of-the-art (SOTA) multi-modal LLMs (MLLMs) (Liu et al., 2024b; Li et al., 2023b) increasingly build on open-source LLM backbones (Touvron et al., 2023; Peng et al., 2023; Jiang et al., 2023), aligning visual features to the LLM embedding space (Li et al., 2023a; Fu et al., 2023; Wu & Xie, 2024; Xu et al., 2024; Shao et al., 2024; Li et al., 2022; 2025; Brown et al., 2020; Achiam et al., 2023; Bai et al., 2023a; Jaech et al., 2024; Zhang et al., 2025; 2024). These models have progressively achieved competitive results in complex tasks requiring high-level perception and reasoning (Li et al., 2024a; Gemini, 2023; Fu et al., 2023; OpenAI, 2023), including spatial reasoning (Chen et al., 2024; Cai et al., 2024), character recognition (Mori et al., 1999), scene understanding (Cordts et al., 2016; Wang et al., 2023; Li et al., 2023c; Chen et al., 2017), action recognition (Jhuang et al., 2013; Herath et al., 2017), and action prediction (Lan et al., 2014; Kong & Fu, 2022), in some cases approaching human-level performance.
Building on these advancements, we conduct a systematic evaluation of vision–language models (VLMs) spanning three distinct origins and capacity scales. To ensure comparability, all models are assessed on a zero-shot image–text reasoning task in a generative setting. The complete list of evaluated models, along with curated model size statistics, is presented in the results section (Figure 3). For analysis, we group the models into the following categories:
@@ 2.3 Examined Vision Language Models
    Open-source VLMs with Multi-Image Reasoning: Includes models with different sizes and other variants such as CogVLM Series (Hong et al., 2024), Qwen series(Qwen-VL (Bai et al., 2023b), Qwen-2 (Wang et al., 2024)), and Blip2 (Li et al., 2023b), LLaVA-Next (Liu et al., 2024a) , which are capable of reasoning over interleaved multiple images and texts.
@@
In total, we processed 37 models for evaluation. All the model performances in intentionality understanding and perspective-taking, together with human baseline performances, are presented here (Figure 2). In order to analyze the reasoning abilities of VLMs, we ask the models to explain their answers after they have given the answers.
@@ 3.1 Overall Performance
Figure 4: VLMs perform significantly better in intentionality understanding compared to perspective-taking. Paired samples t-test: $p=2.62\times 10^{-19}$, $t=17.651$ .
@@ 3.2 Relationship Between Model Performance and Model Size
To quantitatively assess the impact of model size on performance, we conducted a linear regression analysis of accuracy scores against model size across our 37 VLMs. For intentionality understanding, the coefficient of determination is $R^{2}=0.2797$, with a slope of $0.0599$ and an offset of $0.3925$. This positive slope indicates that as model size increases, intentionality understanding improves, which aligns with the scaling law hypothesis—the expectation that larger models generally exhibit better performance on cognitive tasks. The results for perspective-taking reveal a strikingly different trend. The coefficient of determination is much lower at $R^{2}=0.0176$, with a slope of 0.0057-0.0057 and an offset of $0.1437$. The negative slope suggests that, contrary to expectations, larger models do not show significant improvements in perspective-taking ability. In fact, their performance remains largely unchanged or even slightly decreases as they scale. This finding directly contradicts the scaling law hypothesis and suggests that perspective-taking may require fundamentally different cognitive processes that are not naturally enhanced through increased model size alone. This dissociation between intentionality understanding and perspective-taking as models scale highlights a potential limitation in current VLMs. While intentionality understanding appears to benefit from increased parameters and training data, perspective-taking does not seem to follow the same trajectory.